OpenStudy (anonymous):

write a ratio and a percent for 5/16 plz help urgant

OpenStudy (anonymous):

To find ratio\[ratio~means~\huge{:}\]\[\frac{ 5 }{ 16 }=\huge{5:16}\]

OpenStudy (anonymous):

Percentages (%) are used to find 100ths of a quantity.

OpenStudy (anonymous):

To find percentages\[=\frac{ 5 }{ 16 }\times100\]\[=\frac{ 5\times100 }{ 16 }\]\[=\frac{ 500 }{ 16 }\]Then,convert it into decimal form\[=?\]solve it @dracolyteking